# SAM 2++: Tracking Anything at Any Granularity

### v3

@@ Abstract
Video tracking aims at finding the specific target in subsequent frames given its initial state. Due to the varying granularity of target states across different tasks, most existing trackers are tailored to a single task and heavily rely on custom-designed modules within the individual task, which limits their generalization and leads to redundancy in both model design and parameters. To unify video tracking tasks, we present SAM 2++, a unified model towards tracking at any granularity, including masks, boxes, and points. First, to extend target granularity, we design task-specific prompts to encode various task inputs into general prompt embeddings, and a unified decoder to unify diverse task results into a unified form pre-output. Next, to satisfy memory matching, the core operation of tracking, we introduce a task-adaptive memory mechanism that unifies memory across different granularities. Finally, we introduce a customized data engine to support tracking training at any granularity, producing a large and diverse video tracking dataset with rich annotations at three granularities, termed Tracking-Any-Granularity, which represents a comprehensive resource for training and benchmarking on unified tracking. Comprehensive experiments on multiple benchmarks confirm that SAM 2++ sets a new state of the art across diverse tracking tasks at different granularities, establishing a unified and robust tracking framework.
  ‡Equal contribution.  †Corresponding author (lmwang@nju.edu.cn).
@@ 1 Introduction
Video tracking has been a fundamental task in computer vision for decades, aiming to estimate the state of an arbitrary target in video sequences given its initial status. Despite sharing this core objective, the tracking domain has fragmented into several independent sub-tasks based on different target granularities, including Single Object Tracking [lasot, TrackingNet, got10k] (SOT) with bounding box, Video Object Segmentation [davis17, youtube-vos, LVOS_V1] (VOS) with precise pixel-level mask, and Point Tracking [badja, zheng2023pointodyssey, tapvid] with tiny points. This fragmentation based on state granularity has led most video tracking research to focus on a specific task and propose specialized designs only for that task. While this design trend enhances tracking performance, it limits the generalization ability of tracking models across multiple tasks and results in redundancy in both model design and parameters. To unify tasks, current unified vision models typically share feature extraction backbones while employing task-specific branches [zhu2022uniperceiver], convert those tasks into a seq2seq framework [chen2021pix2seq], or share one appearance model for either propagation or association [wang2021UniTrack, yu2024unifiedtt, Unicorn, UNINEXT, OmniTracker]. However, they choose to provide different interfaces for different tasks, rather than seeking a unified visual representation of tracking targets, and ignore the point tracking task.
@@
Unlike them, we observe that these seemingly disparate tracking paradigms fundamentally differ primarily in their state granularity, while sharing the memory matching strategy: the model encodes the previous state into memory, and matches the current features with the stored memory when a new frame is received. Based on this strategy, we decide to unify target states at three different granularities through a uniform memory representation. Recently, Segment Anything Model 2 [ravi2024sam2], a strong foundational model, has been proposed for high-quality video object segmentation given various prompts. Due to its flexible prompt mechanism and powerful mask tracking capabilities, we extend this model to track arbitrary granularity, termed as SAM 2++.
@@
Figure 1: The overall of SAM 2++, including (a) tracking any granularity task, (b) our unified tracking foundation model, and (c) our Tracking-Any-Granularity dataset collected through our data engine. SAM 2++ is capable of tracking targets at any granularity.
@@
Our work includes a model and a dataset (see Fig. 1). To ensure generalized tracking at different granularities, we start by designing task-specific prompts and a unified decoder. Specifically, we introduce corresponding prompts for different tasks in various granularities to encode various task inputs into general prompt embeddings. As for the diverse task output, our unified decoder, which is extended from the Mask Decoder of SAM 2, unifies diverse task results into a unified form pre-output. Next, we found that a simple full parameter-shared approach for task mixing training leads to performance degradation on all tasks, due to the different memory requirements of different tracking tasks. To address this, we introduce a task-adaptive memory mechanism, which adjusts memory representations in response to the unique requirements of each task. This mechanism not only helps to offset the adverse effects of full parameter sharing on the memory mechanism but also achieves mutual promotion among multiple tasks. Finally, to enable ”tracking granularity” capabilities in video, we utilize a data engine to construct a large and diverse video tracking dataset, termed Tracking-Any-Granularity (TAG). The data engine produces training data through an interactive process, where annotators manually label data at varying intervals in different phases. Subsequently, after training on the datasets at different phases, the model is used to annotate the remaining frames, achieving efficient and accurate expansion of the dataset. Unlike most existing video tracking datasets, our dataset provides high-quality annotations at three granularities, including segmentation masks, bounding boxes, and key points, resulting in a vital resource for training and benchmarking unified tracking models. Extensive experiments on several benchmarks from various tasks demonstrate that SAM 2++ enables tracking targets at any granularity with a unified model architecture and consistently outperforms task-specific models in all three tasks.
The main contributions are summarized as follows:
@@
  • We propose a unified framework, termed SAM 2++, towards tracking targets at any granularity by task-specific prompts, a unified decoder, and a task-adaptive memory mechanism for various granularities.
@@
  • We build a data engine that produces training data through an interactive process, resulting in a new large-scale object tracking dataset, Tracking-Any-Granularity (TAG), with high-quality annotations in various granularities.
@@ 2 Related work
Segment Anything Model. SAM [kirillov2023seganysam1] is a foundational model for high-quality segmentation given various prompts, and SAM 2 [ravi2024sam2] extends it to video with streaming memory, effectively handling motion and occlusion, and they inspire many variant models In the image domain, HQ-SAM [sam_hq] enhances segmentation quality through a High-Quality Token, SAMRefiner [lin2025samrefiner] improves fine-grained details via a noise-tolerant prompt, CAT-SAM [xiao2024cat] adopts a conditional tuning approach to adapt to specialized image domains, and SAM-Adapter [chen2023samadapter] incorporates lightweight adapters for improved downstream performance. In the video domain, SAM2Long [ding2024sam2long] employs constrained tree search to reduce error accumulation, SAMURAI [yang2024samurai] uses the Kalman filter to select motion-aware memory, while DAM4SAM [videnovic2024distractor] introduces a distractor-aware memory. SAMWISE [cuttano2024samwise] and AL-Ref-SAM-2 [huang2025unleashing] add additional prompts for more referring tasks. Despite these advances, these works remain task-specific, lacking cross-domain generalization and requiring separate implementations for each application.
@@
Unified Vision Models. Recent years have witnessed significant progress in developing unified vision models that handle multiple tasks through shared architectures and demonstrate strong generalizability and flexibility. Pix2Seq [chen2021pix2seq] reformulates vision tasks as sequence generation problems, Uni-Perceiver [zhu2022uniperceiver] establishes unified representation spaces across modalities with shared encoders and decoders. UniTrack [wang2021UniTrack] demonstrates that video tracking tasks can be solved by a single appearance model with task-specific heads, while Unicorn [Unicorn] and UNINEXT [UNINEXT] unify various tracking paradigms through common frameworks with different representations. Despite their impressive capabilities, these unified approaches predominantly focus on object-level tasks while neglecting finer-grained tasks such as point tracking. Furthermore, they do not take into account unifying video tracking tasks with various granularities through a unified visual representation.
3 Preliminaries: Segment Anything Model 2
@@
The Segment Anything Model (SAM) [kirillov2023seganysam1] is a milestone vision foundation model for class-agnostic image segmentation. It flexibly handles various prompts (box, point, mask) by encoding them into a unified embedding and has established an iterative data engine with model-assisted labeling to address dataset limitations. SAM 2 [ravi2024sam2] extends SAM to promptable video segmentation by introducing a streaming memory that stores previous target information and predictions. It comprises four main components: (i) a hierarchical image encoder that encodes each frame ${I}_{img}$ into image embeddings $F_{img}$, (ii) a prompt encoder, (iii) a memory mechanism (memory encoder, memory bank, memory attention), and (iv) a mask decoder for prediction.
@@
Mask Decoder. The mask decoder takes prompt embedding $\mathcal{P}_{sparse}$ and $\mathcal{P}_{dense}$, memory-conditioned image embeddings $\bar{F}_{img}\in\mathbb{R}^{C/4\times H/16\times W/16}$ (which we will explain latter), and a set of learnable tokens $\mathcal{E}_{tokens}$ as inputs. The learnable tokens contain an existence token $\varepsilon_{obj}\in\mathbb{R}^{C}$ to predict whether the target exists, an IoU token $\varepsilon_{iou}\in\mathbb{R}^{C}$ to predict the result accuracy, and multiple mask tokens $\varepsilon_{mask}^{N}\in\mathbb{R}^{N\times C}$ used to obtain $N$ mask candidates. To fuse the prompt embedding, a Two-Way Transformer $\mathbf{twTrans}$ [ravi2024sam2] processes them as:
@@ 4.1 Unified Task Input and Output Processing
Input Unification via Task-Specific Prompt. Due to the input of the three tracking tasks having inconsistent granularity, we first unify inputs with task-specific prompts for different tasks. The video object segmentation task still adopts mask input ${I}_{mask}$ as its prompt in mask form, and the single object tracking task takes its box input ${I}_{box}$ as the prompt. As for the point tracking task, expect the point coordinates ${I}_{point}$, we add a dense mask ${G}_{point}$ for additional prompt as a Gaussian map centred on the point and parameterised by sigma $\sigma$ and radius $r$ as: $G_{point}=\exp\left(-\frac{\|p-p_{0}\|^{2}}{2\sigma^{2}}\right)\cdot\mathbf{1}_{\{\|p-p_{0}\|\leq r\}}$ to highlights the point in mask form, maintaining consistency with output from Unified Decoder and source for Memory Encoder, which is better than naive {0, 1} mask. More importantly, we gradually decrease the radius and sigma during training to facilitate smoother convergence and more stable learning.
Output Unification via Unified Decoder. To unify the output of various tasks, we extended the Mask Decoder of SAM 2 as Unified Decoder, which also processes memory-conditioned image embeddings, prompt embeddings, and learnable tokens. For the SOT task, the outer box of the mask output $M^{N}_{box}$ cannot be used as task output because the complexity of the mask reduces the accuracy of the box, which focuses on the center point’s position and target scale. Instead, we add a Corner-based Head [yan2021learning_stark], $\mathbf{CornHead}$, which explicitly optimizes for the accuracy and stability of the bounding box and is widely used in the SOT task, to produce box predictions. As for the PT task, rather than direct point coordinates, we obtained point predictions in terms of mask predictions by soft-argmax operation during training or argmax operation during inference. This specific design aims for the output of the point task to be consistent with the source of memory, thereby achieving a unified encoding mask output for memory information of different granularities, and also helps to optimize model training. In summary, our unified decoder can be written as:
@@
where $M_{mask}^{i}$, $B_{box}^{i}$, and $P_{point}^{i}$ are task predictions. represent the $i_{th}$ candidate prediction for three tasks, which are rated by their corresponding iou scores $O_{IoU}^{i}$.
@@ 4.2 Task-adaptive Memory
Tracking models fundamentally localize targets according to their past states, which requires efficient storage and retrieval ability using a memory-matching paradigm: the model first encodes the previous states into memory, then matches current features with memory to accurately represent the target when processing a new frame. Following this paradigm, our model converts mask outputs into memory with Memory Encoder, then applies cross-attention in Memory Attention to match the current frame feature with the feature stored in the memory bank as:
@@
Figure 3: Annotation pipeline of our Tracking-Any-Granularity dataset.
Figure 4: Examples of Tracking-Any-Granularity Dataset.
Therefore, we propose a task-adaptive memory mechanism, which relaxes the uniformity by decoupling only the memory components: each task has its own convolutional Memory Encoder, and each applies an independent LoRA [DBLP:conf/iclr/HuSWALWWC22/lora] in the transformer-based Memory Attention. This decoupled design effectively meets the diverse needs of different tasks and avoids the performance drop seen in a fully parameter-shared model, while keeping the overall structure consistent, enhancing the multi-task processing capability of the model. Since only a small number of parameters are decoupled, the increase in parameter count is minimal. Notably, experiments show that this design enables multiple tasks to promote each other.

### v4

@@ Abstract
Video tracking aims at finding the specific target in subsequent frames given its diverse initial states. Due to the varying granularity of target states across different tasks, most existing trackers are tailored to a single task and rely heavily on custom-designed modules. This specificity limits their generalization, preventing them from effectively utilizing multi-task training data and leading to redundancy in both model design and parameters. Although recent unified vision models share partial architectures across tasks, they usually retain task-specific interfaces and overlook the common tracking principle behind different granularities, leaving a gap for truly unified video tracking. To unify video tracking tasks, we present SAM 2++, a unified framework that can handle target states at different granularities, including masks, boxes, and points, through an integrated design of prompt encoding, output decoding, and memory representation. First, to handle different target granularities, we design task-specific prompts that map diverse task inputs into general prompt embeddings, together with a Unified Decoder that produces task results in a common output form without redesigning the overall pipeline. Next, to satisfy memory matching, the core operation of tracking, we introduce a task-adaptive memory mechanism that unifies memory across different granularities while preserving their distinct state semantics, preventing full parameter sharing from causing interference across granularities. Finally, we introduce Tracking-Any-Granularity, the first large and diverse video tracking dataset with rich annotations at three granularities. It is constructed through a customized data engine with phased manual annotation and model-assisted completion, providing a comprehensive resource for training, benchmarking, and analyzing unified tracking models. With the help of our Tracking-Any-Granularity dataset, comprehensive experiments on multiple benchmarks confirm that SAM 2++ sets a new state of the art across diverse tracking tasks at different granularities, establishing a unified and robust tracking framework. Code, trained models, and dataset are available at https://github.com/MCG-NJU/SAM2-Plus.
Index Terms:
Video tracking, Foundation model, Tracking dataset
@@ 1 Introduction
Video tracking has been a fundamental task in computer vision for decades, aiming to estimate the state of an arbitrary target in video sequences given its initial status. Despite sharing this core objective, the tracking domain has fragmented into several independent sub-tasks based on different target granularities, including Single Object Tracking [25, 51, 34] (SOT) with bounding box, Video Object Segmentation [56, 73, 31] (VOS) with precise pixel-level mask, and Point Tracking [5, 86, 23] with tiny points. All three tasks follow user-specified targets over time, but represent target states as coarse object locations, dense object shapes, or fine-grained keypoints at different granularities. This fragmentation based on state granularity has led most video tracking research to focus on a specific task and propose specialized designs only for that task. While this design trend enhances tracking performance, it limits the generalization ability of tracking and training across multiple tasks and results in redundancy in both model design and parameters. To unify tasks, current unified vision models typically share feature extraction backbones while employing task-specific branches [89], convert those tasks into a seq2seq framework [9], or share one appearance model for either propagation or association [66, 82, 75, 76, 62]. Still, such unification often remains at shared components rather than a common representation of tracking states. However, they choose to provide different interfaces for different tasks, rather than seeking a unified visual representation of tracking targets, and ignore the point tracking task.
@@
Unlike them, we observe that these seemingly disparate tracking paradigms fundamentally differ primarily in their state granularity, while sharing the memory matching strategy, where the model encodes the previous state into memory and matches the current features with the stored memory when a new frame is received. This shared process offers a natural entry point for bridging different target representations. Based on this strategy, we decide to unify target states at three different granularities through a uniform memory representation. Recently, Segment Anything Model 2 [58], a strong foundational model, has been proposed for high-quality video object segmentation given various prompts. Due to its flexible prompt mechanism and powerful mask tracking capabilities, we extend this model to track arbitrary granularity, termed as SAM 2++.
@@
Figure 1: The overall framework of SAM 2++, including (a) tracking at any granularity task, (b) our unified tracking framework, and (c) our Tracking-Any-Granularity dataset collected through our data engine. SAM 2++ is capable of tracking at any granularity.
@@
Our work includes a model and a dataset (see Fig. 1). To ensure generalized tracking at different granularities, we start by designing task-specific prompts and a unified decoder. Specifically, we introduce task-specific prompts to encode different task inputs into general prompt embeddings. As for the diverse task output, our unified decoder, which is extended from the Mask Decoder of SAM 2, unifies diverse task results into a unified form pre-output. With this design, task inputs remain compatible with the SAM 2 prompt interface, while task outputs are converted into representations for the shared tracking pipeline. Next, we found that full parameter sharing in task-mixed training leads to performance degradation on all tasks, due to the different memory requirements of different tracking tasks. To address this, we introduce a task-adaptive memory mechanism, which adjusts memory representations in response to the unique requirements of each task. This mechanism not only helps to offset the adverse effects of full parameter sharing on the memory mechanism but also achieves mutual promotion among multiple tasks. Instead of forcing all tasks to write identical memory features, the task-adaptive memory keeps the shared tracking process while allowing each granularity to maintain its own state semantics.
To support tracking at any granularity in videos, we further construct a large and diverse video tracking dataset, termed Tracking-Any-Granularity (TAG), through a data engine. The data engine produces training data through an interactive process, where annotators manually label data at varying intervals in different phases. Then, after phased training, the model is used to annotate the remaining frames, achieving efficient and accurate expansion of the dataset. Unlike most existing video tracking datasets, our dataset provides high-quality annotations at three granularities, including segmentation masks, bounding boxes, and key points, resulting in a vital resource for training and benchmarking unified tracking models. These annotations make TAG suitable for both mixed training and cross-granularity evaluation. Extensive experiments on several benchmarks across tasks demonstrate that SAM 2++ enables tracking at any granularity with a unified model architecture and consistently outperforms task-specific models in all three tasks. The main contributions are summarized as follows:
@@
  • We propose a unified framework, termed SAM 2++, towards tracking targets at any granularity by task-specific prompts, a unified decoder, and a task-adaptive memory mechanism for various granularities, with a shared memory representation.
@@
  • We build a data engine that produces training data through an interactive process both manually and automatically, resulting in a new large-scale object tracking dataset, Tracking-Any-Granularity (TAG), with high-quality annotations in various granularities for unified training and evaluation.
@@ 2 Related work
Segment Anything Model. SAM [41] is a foundational model for high-quality segmentation given various prompts, and SAM 2 [58] extends it to video with streaming memory, effectively handling motion and occlusion, and they inspire many variant models. In the image domain, HQ-SAM [39] and SAMRefiner [48] improve mask quality and fine-grained details, CAT-SAM [70] and SAM-Adapter [8] adapt SAM to specialized domains, while FastSAM [85], MobileSAM [83], and EfficientSAM [71] reduce its computational cost. Grounded SAM [59] and Semantic-SAM [42] extend SAM-style segmentation toward open-world and granularity-aware perception. In the video domain, SAM2Long [22] employs constrained tree search to reduce error accumulation, SAMURAI [78] uses the Kalman filter to select motion-aware memory, while DAM4SAM [61] introduces a distractor-aware memory. SAM-PT [57] combines SAM with point tracking for point-centric video segmentation, EfficientTAM [72] and EdgeTAM [88] improve the efficiency of SAM 2-style track-anything models, and Segment Any Motion [35] explores motion-aware video segmentation. SAMWISE [19] and AL-Ref-SAM-2 [36] add additional prompts for more referring tasks. Despite these innovations addressing specific challenges across the segmentation and tracking spectrum, they remain largely isolated solutions optimized for individual tasks, overlooking potential synergies between related vision problems and requiring separate specialized implementations for each application domain.
@@
Unified Vision Models. Recent years have witnessed significant progress in developing unified vision models that handle multiple tasks through shared architectures and demonstrate strong generalizability and flexibility. Pix2Seq [9, 10] reformulates vision tasks as sequence generation problems, while Uni-Perceiver [89, 43] establishes unified representation spaces across modalities with shared encoders and decoders. Unified-IO [49] and Painter [64] further explore generalist models with unified task interfaces, and SegGPT [65] studies in-context segmentation across different segmentation forms. UniTrack [66] demonstrates that video tracking tasks can be solved by a single appearance model with task-specific heads, while Unicorn [75], UNINEXT [76], MITS [74], and OmniTracker [62] unify different object-level tracking or segmentation paradigms. Despite their impressive capabilities, these unified approaches predominantly focus on object-level tasks while neglecting finer-grained tasks such as point tracking. Furthermore, most require joint training from scratch on multiple large-scale datasets, demanding substantial computational resources that limit their accessibility and development. Moreover, they do not take into account unifying tracking tasks with various granularities through a unified representation.
Tracking Datasets Existing video tracking datasets are usually constructed for a specific target granularity. VOS datasets such as DAVIS-2017 [56], YoutubeVOS [73], LVOS [31, 32], and MOSE [21] mainly provide mask annotations, with motivations ranging from precise labels and large-scale coverage to long-term tracking and complex scenarios. SOT datasets such as LaSOT [25], GOT-10k [34], TrackingNet [51], and others focus on box trajectories for object-level localization, covering challenges such as large-scale training, long-term tracking, high-frame-rate videos, UAV scenes, language-guided tracking, and abundant object categories. Point tracking datasets such as Perception Test [52], PointOdyssey [86], and TAP-Vid [23] provide keypoint or arbitrary-point trajectories for fine-grained evaluation, including multi-modal, long-term, real-world, and arbitrary-point settings. Although these datasets have advanced their respective tasks, they are typically annotated for a single task and a single target-state granularity. In contrast, our Tracking-Any-Granularity (TAG) dataset provides segmentation masks, bounding boxes, and key points for the same video tracking data, making it the only dataset in our comparison that supports all three granularities simultaneously for unified tracking training and evaluation.
3 Preliminaries: Segment Anything Model
@@
The Segment Anything Model (SAM) [41] is a milestone vision foundation model for class-agnostic image segmentation. It flexibly handles various prompts (box, point, mask) by encoding them into a unified embedding and has established an iterative data engine with model-assisted labeling to address dataset limitations. SAM 2 [58] extends SAM to promptable video segmentation by introducing a streaming memory that stores previous target information and predictions. It comprises four main components: (i) a hierarchical image encoder that encodes each frame ${I}_{img}$ into image embeddings $F_{img}$, (ii) a prompt encoder, (iii) a memory mechanism (memory encoder, memory bank, memory attention), and (iv) a mask decoder for prediction.
@@
Mask Decoder. The mask decoder takes prompt embedding $\mathcal{P}_{sparse}$ and $\mathcal{P}_{dense}$, memory-conditioned image embeddings $\bar{F}_{img}\in\mathbb{R}^{C/4\times H/16\times W/16}$ (which we will explain later), and a set of learnable tokens $\mathcal{E}_{tokens}$ as inputs. The learnable tokens contain an existence token $\varepsilon_{obj}\in\mathbb{R}^{C}$ to predict whether the target exists, an IoU token $\varepsilon_{iou}\in\mathbb{R}^{C}$ to predict the result accuracy, and multiple mask tokens $\varepsilon_{mask}^{N}\in\mathbb{R}^{N\times C}$ used to obtain $N$ mask candidates. To fuse the prompt embedding, a Two-Way Transformer $\mathbf{twTrans}$ [58] processes them as:
@@ 4.1 Unified Task Input and Output Processing
Input Unification via Task-Specific Prompt. Due to the input of the three tracking tasks having inconsistent granularity, we first unify inputs with task-specific prompts for different tasks, while keeping the original Prompt Encoder unchanged. This preserves the prompt space of SAM 2 and avoids introducing extra input encoders for individual tracking tasks. Mask tracking adopts the {0, 1} mask ${I}_{mask}$ as the prompt to accurately describe the target shape and boundary, while box tracking uses the bounding box ${I}_{box}$ as the prompt, represented by its top-left and bottom-right corners. Point tracking uses the point coordinate ${I}_{point}$ together with an additional Gaussian mask ${G}_{point}$, which is centred on the point and parameterised by sigma $\sigma$ and radius $r$ as: $G_{point}=\exp\left(-\frac{\|p-p_{0}\|^{2}}{2\sigma^{2}}\right)\cdot\mathbf{1}_{\{\|p-p_{0}\|\leq r\}}$. This Gaussian mask highlights the point in mask form, maintains consistency with the output from Unified Decoder and the source for Memory Encoder, and better represents the target than a naive {0, 1} mask. More importantly, we gradually decrease the radius and sigma during training to facilitate smoother convergence and more stable learning.
Output Unification via Unified Decoder. To unify the output of various tasks, we extend the Mask Decoder of SAM 2 as Unified Decoder, which also processes memory-conditioned image embeddings, prompt embeddings, and learnable tokens, and make lightweight task-specific changes on top of the shared mask-form output. For the SOT task, directly deriving the outer box from the mask output $M^{N}_{box}$ cannot provide an ideal task output, because the complexity of the mask reduces box accuracy and the outer-box operation lacks a direct training gradient. Instead, we add a Corner-based Head [77], $\mathbf{CornHead}$, which explicitly optimizes the position and scale of the bounding box and is widely used in the SOT task, to produce box predictions. As for the PT task, rather than adding an extra coordinate regression head, we obtain point coordinates from the mask output through a soft-argmax operation during training to ensure gradient flow and an argmax operation during inference. This design aligns point prediction with the Gaussian prompt and the mask source used by Memory Encoder, thereby keeping the point output consistent with the unified memory representation and helping optimize model training. In summary, our unified decoder can be written as:
@@
where $M_{mask}^{i}$, $B_{box}^{i}$, and $P_{point}^{i}$ are predictions. represent the $i_{th}$ candidate prediction for three tasks, which are rated by their corresponding iou scores $O_{IoU}^{i}$.
@@ 4.2 Task-adaptive Memory
Tracking models fundamentally localize targets according to their past states, which requires efficient storage and retrieval ability using a memory-matching paradigm: the model first encodes the previous states into memory, then matches current features with memory to accurately represent the target when processing a new frame. Following this paradigm, our model converts mask outputs into memory with Memory Encoder, then applies cross-attention in Memory Attention to match the current frame feature with the feature stored in the memory bank as
@@
Therefore, we propose a task-adaptive memory mechanism, which relaxes the uniformity by decoupling only the memory components. Specifically, for the convolutional Memory Encoder, each task has its own copy; for the transformer-based Memory Attention, each task is equipped with an independent LoRA [33]. This decoupled design effectively meets the diverse needs of different tasks and avoids the performance drop seen in a fully parameter-shared model, while keeping the overall structure consistent and introducing only a minimal parameter increase. Notably, experiments show that this design enables multiple tasks to promote each other. Ultimately, we achieve unified multi-task processing with a largely shared memory representation.